### v1

# Status Report on the Chicago-Carnegie Hubble Program (CCHP): Three Independent Astrophysical Determinations of the Hubble Constant Using the James Webb Space Telescope

Abstract
We present the latest results from the Chicago Carnegie Hubble Program ( CCHP) to measure the Hubble constant using data from the James Webb Space Telescope (JWST). This program is based upon three independent methods: (1) Tip of the Red Giant Branch (TRGB) stars, (2) JAGB (J-Region Asymptotic Giant Branch) stars, and (3) Cepheids. Our program includes 10 nearby galaxies, each hosting Type Ia supernovae (SNe Ia), suitable for measuring the Hubble constant (). It also includes NGC 4258, which has a geometric distance, setting the zero point for all three methods. The JWST observations have significantly higher signal-to-noise and finer angular resolution than previous observations with the Hubble Space Telescope (HST). We find three independent values of = 69.85 1.75 (stat) 1.54 (sys) for the TRGB, = 67.96 1.85 (stat) 1.90 (sys) km s-1 Mpc-1 for the JAGB, and = 72.05 1.86 (stat) 3.10 (sys) for Cepheids. Tying into SNe Ia, and combining these methods adopting a flat prior, yields our current estimate of = 69.96 1.05 (stat) 1.12 (sys) km s-1 Mpc-1. The distances measured using the TRGB and the JAGB method agree at the 1% level, but differ from the Cepheid distances at the 2.5-4% level. The value of based on these two methods with JWST data alone is = 69.03 1.75 (total error) km s-1 Mpc-1. These numbers are consistent with the current standard CDM model, without the need for the inclusion of additional new physics. Future JWST data will be required to increase the precision and accuracy of the local distance scale.
1 Introduction
The year 2024 marks just over a century since Edwin Hubble’s famous discovery of a single Cepheid variable in the Andromeda galaxy222The discovery plate was obtained 6 October 1923, on which Hubble wrote VAR! in red ink. See https://obs.carnegiescience.edu/PAST/m31var.. Hubble’s subsequent measurements of extragalactic distances (Hubble, 1929) were based, in part, on the Cepheid period-luminosity (PL) relation (also now widely known as the Leavitt Law, Leavitt (1908)). For 70 years a number of unrecognized challenges (e.g., reddening and dimming due to the effects of interstellar dust, errors in photometric zero points, effects due to differing metal abundances, crowding/blending as a result of insufficient resolution, and the inclusion of some (secondary) distance indicators) turned out to have large systematic effects, and combined together to make it virtually impossible for the Hubble consant () to be measured from the ground to better than a factor of two uncertainty.
This impasse was largely overcome by technological advances: (1) the widespread availability of two-dimensional, linear detectors operating in the optical and near-infrared, beginning in the 1980s (e.g. McGonegal et al., 1982; Freedman et al., 1985, 1991), and (2) ultimately the launch of the Hubble Space Telescope (HST), and undertaking of the Hubble Key Project (Freedman et al.). These order-of-magnitude advances along the three orthogonal axes of sensitivity, wavelength coverage and angular resolution made it possible to reduce the 100% (‘factor-of-two’) uncertainty on the Hubble constant down to 10%, yielding = 72 3 (stat) 7 [sys] (Freedman et al., 2001). Over the following decades this space-based measurement of has been confirmed by multiple subsequent analyses (e.g., Riess et al., 2009; Freedman et al., 2012; Riess et al., 2016, 2022), all based on a Cepheid calibration of distant SNe Ia, subsequently taken into the more distant Hubble flow.
Twenty-five years later, one of the outstanding questions in cosmology is this: Have we overcome the systematic effects (both known and perhaps still unknown) that can affect measurements of the astrophysical distance scale at a sufficient level to require that our current standard cosmological model (CDM) is now in need of additional physics? This question has arisen with the emergence of a completely new method for inferring the value of that is totally independent of the Cepheid calibration, and is, instead, based upon modeling measurements of the fluctuations in the cosmic microwave background. Under the assumption of the standard (CDM) cosmological model, the CMB measurements from the Planck satellite predict a current expansion rate of 67.4 0.5 km s-1 Mpc-1 (i.e., with better than 1% precision)(Planck Collaboration et al., 2020). Consistent results are also obtained from the Atacama Cosmology Telescope (ACT, (e.g., Madhavacheril et al., 2024)), and from the South Pole Telescope (SPT, (e.g., Balkenhol et al., 2023)). However, the CMB results differ from the most recent (local) measurements of based on distant Type Ia supernovae (SNe Ia) in the unperturbed Hubble flow, calibrated using HST observations of Cepheids. The difference is at a level of 5, as estimated by (Riess et al., 2022, hereafter R22), a discrepancy known as the Hubble tension. If this level of discrepancy is confirmed, it suggests the existence of new physics (particles or fields) not yet constrained by the standard (CDM) cosmological model. To date, however, no plausible changes or additions have emerged that allow for values of as high as 73 km s-1 Mpc-1 (see Di Valentino et al., 2021).
Consistent with CMB measurements, lower values of are also attained from measurements of baryon acoustic oscillations (BAO), or fluctuations in the matter density. Most recently the Dark Energy Spectroscopic Instrument (DESI) Collaboration Data Release 1 (DR1) (DESI Collaboration et al., 2024) find = 68.52 0.62 km s-1 Mpc-1. This value was obtained by calibrating the standard BAO ruler based on a prior for the baryon density from Big Bang Nucleosynthesis (BBN), in addition to a measurement of the CMB acoustic angular scale, . (The CMB acoustic angular scale has been measured to a very high precision of 0.03% (Planck Collaboration et al., 2020)). A slightly less precise (but nonetheless the same) result is found using the BBN calibration alone, giving = 68.53 0.80 km s-1 Mpc-1. This result is completely independent of the CMB measurements. If instead, a calibration based on CMB measurement of the sound horizon distance, rd is used, the resulting value of is 69.29 0.87 km s-1 Mpc-1.
HST has now served for more than 30 years as the primary, space-based instrument for the discovery and measurement of extragalactic Cepheid variables, and the local determination of (e.g., the Hubble Key Project (Freedman et al., 2001; Riess et al., 2022)). With its high resolution and sensitivity at optical wavelengths (where the amplitudes of Cepheids are large), HST is an ideal instrument for discovery of extragalactic Cepheids. In contrast, JWST is the best-equipped telescope for improving the accuracy of measurements of , due to its greater sensitivity at long wavelengths and its even higher spatial resolution relative to HST. The long-wavelength capability of JWST also makes it an ideal facility for the study of other distance indicators; for example, the redder TRGB and JAGB (J-Region Asymptotic Giant Branch) populations. Additionally, in the near- and mid-infrared, interstellar extinction is significantly lower than in the optical (e.g., AJ and A[4.4] are smaller by factors of 4 and 20 respectively, relative to the visual extinction, AV; and factors of 2 and 10 lower relative to the I-band extinction AI (Cardelli et al., 1989; Indebetouw et al., 2005). Moreover, the science performance of JWST has exceeded initial estimates of its sensitivity, stability, image quality, as well as spectral range (Rigby et al., 2022). NIRCam (F115W) imaging from JWST (Rieke et al., 2023) has a sampling resolution four times higher than HST (F160W), with a FWHM of 0.04 arcsec on JWST, versus 0.15 arcsec on HST. Of some potential concern, at red wavelengths, red giant and bright asymptotic giant branch stars can impact the photometry of the Cepheids (due to crowding and blending effects), exacerbating these effects in the red, compared to optical wavelengths. Importantly, with four times better resolution than HST at these near-infrared wavelengths, crowding effects are decreased by more than an order of magnitude in flux when using JWST.
In this paper, we present our results from a new long-term Chicago-Carnegie Hubble Program (CCHP) using the James Webb Space Telescope (JWST). The aim of the program is to reduce the systematic uncertainty in the local extragalactic distance scale, and ultimately the measurement of . Our current goals are: (1) to use three independent stellar distance indicators (Cepheids, the TRGB, and JAGB stars) to obtain three high-precision distances to each calibrating galaxy, thereby reducing the overall systematic distance uncertainties, (2) to make use of the high resolution of JWST to understand and reduce the possible effects of crowding and blending of Cepheids in previous HST photometry, (3) to improve the corrections for dust, and (4) to improve constraints on the effects of metallicity on the Cepheid Leavitt law. Preliminary results from this program have been published in (Lee et al., 2024a; Owens+, 2024a, submitted; Hoyt et al., 2024; Freedman & Madore, 2023a, b). Details of the measurements for each of the methods discussed here are presented in three companion papers: (Owens+, 2024b, in prep; Hoyt+, 2024b, submitted; Lee+, 2024, submitted, hereafter O24, H24, L24, respectively). A fourth paper describes the details of the photometric reductions (Jang+, 2024, in prep). We note also that recent JWST results have also been reported by Yuan et al. (2022); Riess et al. (2023, 2024).
The outline of this paper is as follows: In §2 we provide an overview of the CCHP. We then describe the galaxy sample and analysis in §3, our blinding procedure in §4, and the calibration in §5. In §6 we provide a description of the three methods used in this paper: Cepheids, the TRGB and the JAGB. In §7 we summarize our steps for the measurement of our new distances. In §8 we compare the distances obtained for the three methods, followed by a description of the calibration of SNe Ia and determination of in §9. In §10 we discuss the uncertainty that results from the small number of SN Ia calibrating galaxies, and §11 presents a discussion of the overall uncertainties and an error budget. In §12 we compare our results with previously published data. Section 13 provides a description of future prospects. Finally, in §14 we present our conclusions.
2 The Chicago Carnegie Hubble Program: An Overview
The CCHP was designed as a follow-up to the HST Key Project, (originally named the Carnegie Hubble Program, CHP). Its primary goal was to decrease the systematic uncertainties in the measurement of . Initially it began with a program designed to utilize the mid-infrared capability of the Spitzer Space Telescope, in anticipation of the parallax satellite Gaia. Our early infrared initiative provided 3.6 and 4.5 m data for Cepheids in the Milky Way (Monson et al., 2012), the Large (LMC) and Small (SMC) Magellanic Clouds and the Local Group dwarf irregular galaxy, IC 1613 (Scowcroft et al., 2011, 2013, 2016a, 2016b). Based on a 3.6 m distance to the LMC of 18.477 0.033 mag, a recalibration of the HST Key Project data (Freedman et al., 2012) resulted in a value of = 74.3 2.1 (2.8%) km s-1 Mpc-1, reducing the formal uncertainty by a factor of three.
In its second phase, the CCHP undertook a calibration of the Tip of the Red Giant Branch (TRGB) using the HSTAdvanced Camera for Surveys (ACS) to derive distances to nearby galaxies in which Type Ia supernovae (SNe Ia) had been discovered, and thus provided a calibration independent from that based on Cepheids alone. TRGB distances were measured to 15 galaxies that were known hosts to 18 SNe Ia. The halos of these galaxies were intentionally targeted to minimize the effects of dust, as well as to minimize contamination by younger and brighter disk asymptotic giant branch (AGB) stars. At the same time, crowding was also reduced by pre-selecting low surface brightness regions in the outer halo. The resulting distances were then used to tie into a more distant sample of SNe Ia, observed with high cadence and 9 filters (from the ultraviolet to the near infrared) as part of the Carnegie Supernova Project ( CSP) Krisciunas et al. (2017). This effort resulted in a value of = 69.8 0.6 (stat) 1.6 [sys] km s-1 Mpc-1 Freedman et al. (2019, 2020); Freedman (2021). The results were little changed if instead the SNe Ia catalog from the SHoES collaboration Scolnic et al. (2015) was adopted, giving = 70.4 1.4 1.6 km s-1 Mpc-1 Freedman et al. (2019).
In the current and third phase of the CCHP, we have undertaken a three-pronged approach to measuring using Cycle 1 JWST observations (JWST-GO-1995: P.I. W. L. Freedman). We observed 10 nearby galaxies that are hosts to type SNe Ia, and measured three independent distances using Cepheids, the TRGB and JAGB stars. In addition, we obtained observations of NGC 4258, a galaxy that is host to H2O megamasers, and which provides a geometric distance and our absolute calibration. The program was designed to deal specifically with known systematic effects in the measurement of distances to nearby galaxies: (a) extinction and reddening by dust along the total line of sight to these objects, (b) metallicity effects, and (c) crowding/blending of stellar images. Simply getting more nearby galaxy distances (decreasing the statistical uncertainties) is insufficient to confirm or refute whether new physics beyond the standard cosmological model is required. At this time, systematic uncertainties are (and have historically been) the dominant component of the error budget.
3 The Galaxy Sample, Data Acquisition and Analysis
The galaxies chosen for this program have (1) well-observed SNe Ia with well-defined light curves and peak magnitudes, (2) previously-discovered Cepheid variables (Freedman et al., 2001; Riess et al., 2020), and (3) are located at distances 23 Mpc, suitable for measuring accurate Cepheid, TRGB and JAGB (carbon-star) distances. All three of these methods have high precision and can be independently used to calibrate SNe Ia. Images of our 11 target galaxies are shown in Figure 1.
Observations were obtained using the JWST Near-Infrared Camera (NIRCam; Rieke et al. (2023)), with the filter (comparable to ground-based band) and the F356W filter at 3.6m. The observations were carried out over a 13-month period of time from November 2022 to January 2024. The first observations were obtained for the galaxies NGC 7250, NGC 4536 and NGC 3972 (Lee et al. (2024a); Freedman & Madore (2023a, b); Owens+ (2024a, submitted); Hoyt et al. (2024), making use of the F444W filter at 4.4 m, to both provide a metallicity test for Cepheids, and a long color baseline for the discovery and color discrimination of the JAGB stars. We then switched to F356W for the rest of the target sample, owing to its higher sensitivity and better spatial sampling. However, the F444W filter contains a CO bandhead that is sensitive to metallicity (Scowcroft et al., 2016b), and it is being used to carry out a test for metallicity effects in the two nearest galaxies, M101 and NGC 4258 (at distances 7.5 Mpc) where the highest resolution can be achieved.
NIRCam covers a total area of 9.7 arcmin2 field with a 44 arcsec gap between two 2.2 × 2.2 arcmin regions. Our target fields were chosen to optimize the inclusion of (1) the largest possible number of known Cepheids in the inner disk, (2) portions of the extended outer disk to detect carbon stars, and (3) with a rotation angle optimized for the detection of halo red giants (H24).
A detailed description of our data reduction procedures can be found in Jang+ (2024, in prep), hereafter J24. Here we provide a brief overview. The images333We obtained the stage 2 _cal images with jwst_1149.pmap from the MAST archive. were processed primarily using the NIRCam module in the software package DOLPHOT, updated for JWST data analysis (Weisz et al., 2023, 2024). We found a significant improvement in the point-spread modeling for the newer version of DOLPHOT444DOLPHOT 2.0 NIRCAM module sources (updated 4 Feb 2024), resulting in considerably better source identification and image subtraction using a ‘warmstart’ mode. In addition we are carrying out a parallel and independent analysis using DAOPHOT (Stetson, 1987) in order to provide a quantitative constraint on photometric errors that might arise due to differences in point-spread-function fitting and sky-subtraction approaches in crowded fields. In the case of one of our outer fields in the anchor galaxy, NGC 4258, upon which our current calibration is based, we find excellent agreement between the DOLPHOT and DAOPHOT reductions at a level of 0.002 mag (0.1%) with a total of 0.02 mag down to 3 mag below the TRGB. The comparisons for our other galaxies are ongoing. We also tested the automatic aperture corrections (apcorr = 1) provided by DOLPHOT by comparing the overlapping regions of dithered images, as well as by carrying out comparisons with manual/visual determinations of the aperture corrections. These tests confirmed that the automatic aperture corrections were good to a level of 0.02 mag. Finally, all of our initial data analysis was carried out using a blinding procedure, as described in §4 below.
We note that for our Cepheid analysis, we made use of archival F350LP, F555W and F814W HST data in combination with the F115W JWST near-infrared data. For these archival HST F350LP, F555W, and F814W data, we have adopted the same photometric calibration for the Cepheid fields as given by R22. We did not make use of the HST F160W filter, as the near-infrared resolution with NIRCam JWST surpasses that of HST, and the new data are far less affected by crowding (see Figure 2). The three wavelengths (F555W, F814W and F115W) jointly allow a more precise measurement of the extinction. The Cepheid data were reduced using the ‘warmstart’ mode in DOLPHOT where the x,y coordinates of the sources were determined from a stacked F115W-band image. The optical data were then reduced separately using the same source coordinates (see O24 for more details). For the TRGB and JAGB analyses, we used the NIRCam F115W and the F356W and/or F444W data, and did not use any optical HST data. The source coordinates were obtained from the F115W images, and again used in ‘warmstart’ mode for the longer wavelength data (for more details see H24 and L24).
We show in Figure 3 a color-magnitude (F115W versus F115W-F444W) diagram with the locations of all three of the stellar distance indicators indicated. The Cepheids are the brightest objects, followed by the redder JAGB stars and then the TRGB stars. The power of this program can be seen at a glance: the same detector, pixel scale, point spread functions, photometric reduction packages, and calibration are being applied to all three methods, simultaneously. A deliberate placing and orientation of the two NIRCam detectors has allowed the TRGB to be measured in the halo, the Cepheids in the inner disk, and the JAGB stars in the outer disk.
As described in O24, we injected 1,000 artificial stars into the frames, surrounding each of the known Cepheids. Magnitudes of the artificial stars were retrieved in the same manner as for the real stars in the frames. These measurements then provide a mean (statistical) crowding correction. Application of, and the uncertainties in, these statistical corrections are described in O24.Artificial star corrections were not needed for the TRGB or JAGB measurements. The crowding corrections are most significant for the inner disk (Cepheid) fields. As expected, they are significantly less for JWST than for HST. For the NIRCam filter, the average random (statistical) corrections measured from the artificial stars ranged from -0.003 mag (M101) to -0.012 mag (NGC 3972). For the HST filter, the average corrections ranged from -0.010 mag (NGC 4424) to -0.096 mag (NGC 2442); and for the HST filter, the average corrections ranged from -0.033 mag (M101) to -0.133 mag (NGC 2442).
Galaxy | SN Name | Morphological TypeaaAs described in the text, when combining the methods for the final analysis, the uncertainty for the JAGB method was increased by a factor of 1.18 (given by the ratio of the average values for TRGB and Cepheids to that of the lower value for the JAGB), so that the JAGB method does not unduly contribute more weight than the other two methods. | O/H] bbfootnotemark: | |
M101 | SN 2011fe | SAB(rs)cd | 0.10 dex | |
NGC 1365 | SN 2012fr | SB(s)b | -0.14 dex | |
NGC 2442 | SN 2015F | SAB(s)bc pec | 0.00 dex | |
NGC 3972 | SN 2011by | SA(s)bc | 0.03 dex | |
NGC 4038 | SN 2007sr | SB(s)m pec | 0.03 dex | |
NGC 4424 | SN 2012cg | SB(s)a: | 0.06 dex | |
NGC 4536 | SN 1981B | SAB(rs)bc | -0.15 dex | |
NGC 4639 | SN 1990N | SAB(rs)bc | -0.01 dex | |
NGC 5643 | SN 2013aa, SN 2017cbv | SAB(rs)c | 0.13 dex | |
NGC 7250 | SN 2013dy | Sdm: | -0.28 dex |
References. — (a) NED (NASA Extragalactic Database)
4 Our Blinding Procedure
Our initial goal was to undertake all of the data analysis without knowledge of the true zero-point calibration, starting from the receipt of the raw data frames, through to the final analysis and determination of for the three independent methods. In practice, the blinding experiment was (completely) successful only for the JAGB analysis, as described below.
To set up the blinding procedure initially, the frames were processed using DOLPHOT, as described above. Random numbers were then generated and added separately to each of the photometry catalogs based on the raw data frames. The same random offset was added to all passbands in each galaxy, preserving (arbitrary) colors. The blinding was such that during the entire year and a half of the photometric analysis, no one in the group had any knowledge of what the true distances or the value of might be.
One week before unblinding, one team member was given access to the unblinded photometry. This step was taken to test for any potential catastrophic errors somewhere in the blinded analysis. None were found. As part of this effort, an independent analysis, using independent software (fitting of the PL relations, measurement of the TRGB and JAGB luminosity functions) was initiated.
For the TRGB, spatial cuts of the data, selection of TRGB stars, slopes of the TRGB, rectification, Sobel edge detection and measurement of the TRGB were made with the random offsets applied to the individual data frames. Finally, TRGB distances, with arbitrary zero-points, were measured with respect to NGC 4258 (the latter also having a random offset). Similarly, in the case of the JAGB stars, all analysis through spatial cuts (and decisions about whether there was convergence in the spatial cuts), measurement of the luminosity functions, and measurement of distances were carried out with blinded, arbitrary zero points.
In the case of the Cepheid distance scale, and before unblinding, the R22 Cepheids were identified in the frames, light curves were generated, mean magnitudes and PL relations were measured, artificial star experiments were carried out, and mean bias corrections were applied to the PL relations. The artificial star tests were completed for half the sample before the unblinding step. These tests, which can take several weeks per galaxy, continued until completion, and the corrections were added to the unblinded photometry. The corrections are simply additive, involve no choices, and they did not interfere with the blinding approach.
Finally, the entire group met in person on March 13, 2024, and we unblinded the photometry with everyone present, with five of the six co-authors seeing the unblinded results for the first time. We applied the calibration based on NGC 4258 (which until that time had also had an arbitrary zero point) and compared the distances that had been obtained independently using the three different methods (as described in §8 below). At our unblinding meeting, we applied the new local calibrating galaxy distances to obtain distances to the Carnegie Supernova Program and Pantheon+ SN Ia data (§9.3), and independently determined preliminary values of based on the two SN Ia samples. At the time of unblinding the agreement among the three distance indicators was extremely good.
No further updates were made to the JAGB analysis following unblinding. We can consider the JAGB distances obtained in a completely blinded fashion, and this rung of the blinding experiment a success from start to finish. In the case of the TRGB, all aspects of the analysis remained unchanged after unblinding, except for the sample-wide error estimation and the spatial selection for four of the thirteen fields, all of which had the four largest uncertainties in the blinded analysis. Closer inspection of the color-magnitude diagrams in all of these cases revealed an underpopulated tip that was filled out with the post-unblinding larger spatial selection. (For further details see H24.) The sense of the changes was to decrease the distances in these cases, with a subsequent increase to the value of . These differences were within one-sigma of the blinded results. Finally, subsequent to our unblinding meeting, an error was discovered in our Cepheid analysis pipeline, which resulted in a decrease to our preliminary Cepheid distances (and also a subsequent increase in the value of ). These differences resulted in larger changes at a three-sigma level.
5 NGC 4258 and Photometric Zero-point Calibration
Our zero-point calibration for the JWST data is based on the geometric distance to the galaxy, NGC 4258. For ground and HST distance scale measurements, the Large Magellanic Cloud (LMC) and the Milky Way also serve as calibrators. Unfortunately, the Milky Way Cepheids, TRGB and JAGB stars are too bright to be observed with JWST. NGC 4258 is a nearby, highly-inclined spiral galaxy located at a distance of 7.6 Mpc. A geometric distance to the galaxy has been measured via its H2O masers, which are orbiting within an accretion disk inclined at 72∘, surrounding a supermassive black hole (see Humphreys et al., 2013; Reid et al., 2019) and megamaser. The most recent geometric distance to NGC 4258 is 29.397 0.024 (stat) 0.022 [sys] mag (Reid et al., 2019), a 1.5% measurement.
There are several advantages to a calibration based on NGC 4258. Importantly, it can be applied consistently across all three of the methods (Cepheids, TRGB and JAGB stars). Unlike the case of the nearby Milky Way Cepheids and those in the LMC, crowding and blending effects for NGC 4258 are comparable to those of the more distant galaxies, and the range of magnitudes for the Cepheids, TRGB and JAGB stars in both the calibrator and target galaxies allows them to be observed with the same telescope and instrument. Furthermore, by adopting the distance to NGC 4258, it requires simply JWST flight magnitudes alone, without recourse to ground-based tie-ins. Finally, the metallicity of NGC 4258 is comparable to the mean metallicity of the Cepheid fields used in our target galaxies.
Inspecting Table 3, our JWST galaxy sample has a range in average metallicity going from -0.28 [O/H] 0.13 dex, with a mean of -0.023 dex. This value is comparable to that of the R22 sample of 37 galaxies (their Table 3), which has a mean metallicity of -0.037 dex. NGC 4258 has a metallicity of [O/H] = -0.10 dex. The difference in average metallicity for our sample with respect to NGC 4258 is -0.077 dex. For a value of = -0.2 mag/dex (as adopted by R22), this translates to a Cepheid metallicity correction in the mean of less than a 1% (i.e., -0.7%) to .
An external check of the NGC 4258 calibration can be obtained by comparing previous results for the case where the determination of was based on several anchors (NGC 4258, Milky Way, LMC, and SMC) rather than NGC 4258 alone. In the case of the TRGB, applying distances for these four anchors resulted in (internally) good agreement: Freedman (2021) found values of 69.7 1.0 (stat) 2.0 [sys] for NGC 4258, 69.3 0.8 (stat) 3.5 [sys] for the Milky Way, 69.9 0.5 (stat) 1.6 [sys] for the LMC and 69.5 1.0 (stat) 1.7 [sys] for the SMC calibrations, respectively. Similarly, good internal agreement for the Cepheid calibration was also found by R22 with values of 72.51 1.54 for NGC 4258, 73.02 1.19 for the Milky Way, and 73.59 1.36 for the LMC calibrations, respectively. Thus, having NGC 4258 as the sole calibrator cannot explain the reason for the difference in between 69 and 73: the internal agreement amongst the anchors is excellent for both the TRGB and Cepheids.
6 Measurement of Distances
In this section we give a brief overview of the three methods used in this study, and compare their advantages and disadvantages. There is no single perfect method. The weaknesses of each method serve to underscore the importance of having independent methods for constraining overall systematic effects in the local distance scale.
6.1 TRGB
The use of bright red giant branch stars as distance indicators goes back over a century to Harlow Shapley, who measured the brightest stars in globular clusters, as one of his techniques for determining the size of the Milky Way galaxy (Shapley, 1918). Beginning in the 1980s, the availability of CCD detectors with red sensitivity and high quantum efficiency resulted in a resurgence of interest in these stars, this time in their application to the extragalactic distance scale (Mould & Kristian, 1986; Da Costa & Armandroff, 1990; Lee et al., 1993).
The TRGB now provides one of the most precise and accurate means of measuring distances to nearby galaxies, comparable to the Cepheid Leavitt law. In practice, the observed color-magnitude diagrams of old red giant branch stars display a distinct edge/discontinuity in the red giant branch (RGB) luminosity function (LF), corresponding to the core helium-flash luminosity at the end phase of RGB evolution for low-mass stars. Measurement of the TRGB edge has repeatedly been shown to be an excellent standard candle in the I band (Lee et al., 1993; Rizzi et al., 2007; Salaris et al., 2002; Madore et al., 2009; Freedman et al., 2019; Jang et al., 2021), and it is a standardizable candle in the near infrared (Dalcanton et al. (2012); Wu et al. (2014); Madore et al. (2018); Durbin et al. (2020)). Details of the method can be found in a number of reviews (e.g., Madore & Freedman, 1999; Freedman & Madore, 2010; Beaton et al., 2019; Madore & Freedman, 2023).
A strength of the TRGB method is that the underlying theory for why it is an excellent (empirical) standard candle is well understood (Salaris & Cassisi, 1997, 2005; Bildsten et al., 2012; Kippenhahn et al., 2013; Serenelli et al., 2017). After leaving the main sequence, low-mass stars with masses develop a degenerate helium core. They ascend the red giant branch, with their luminosity being powered by a hydrogen-burning shell surrounding the core. The freshly formed helium adds to the core mass until it reaches a value of about 0.5 M⊙ independent of the initial mass of the star. When the temperature reaches about 108K in the isothermal core, it enables the triple-alpha process (helium burning) to commence. Because the degenerate core cannot expand, a thermonuclear runaway ensues (the core helium flash), injecting energy that eventually lifts the electron degeneracy. The star then rapidly evolves away from the red giant branch and descends in luminosity onto the horizontal branch or the red clump, where it undergoes stable core helium burning at a lower-luminosity.
The advantages of the TRGB method are: (1) The discontinuity in the observed TRGB LF is empirically simple to identify and measure. (2) The physical mechanism for the TRGB (the core helium flash) is well understood. (3) If applied in the outer halos of galaxies, where the surface brightness of the galaxy is low, then the overlapping (i.e., crowding or blending effects) of the stellar point spread functions is minimal. (4) In the halos of galaxies, the effects of dust are small (e.g., Ménard et al., 2010). (5) The metallicity of a star on the red giant branch relates directly to its observed color (e.g., Da Costa & Armandroff, 1990; Carretta & Bragaglia, 1998). (6) For a measured TRGB slope in a given passband, the slope of the giant branch luminosity in a different passband is uniquely defined mathematically (Madore & Freedman, 2020).
The disadvantages of the TRGB method are: (1) Care must be taken to ensure that the regions in the target and calibrating galaxies are comparable in terms of surface brightness and line-of-sight column density of dust. (2) Target field placement is critical. Too close in and crowding/blending/reddening become issues. Too far out and the number density of stars is too low for a precise measurement of the tip. (3) Occasionally adjacent peaks of comparable strength can be found in poorly sampled luminosity functions, complicating (or making impossible) an accurate measurement of the true tip. Cases (2) and (3) can be remedied with the acquisition of augmented sample sizes.
6.2 JAGB Stars
Although relatively new in the context of the extragalactic distance scale, J-region AGB (JAGB) stars were, in fact, first identified in the LMC nearly a quarter of a century ago, as a distinct class of objects (Nikolaev & Weinberg, 2000; Weinberg & Nikolaev, 2001), when they were used to determine the back-to-front geometry of the LMC. The JAGB method is now one of the most promising methods for measuring the distances to nearby galaxies. (1) These thermally-pulsating AGB carbon stars have a nearly constant luminosity in the near-infrared J band (at 1.2 m), and (2) they have a low intrinsic dispersion of only 0.2 mag Nikolaev & Weinberg (2000); Madore & Freedman (2020). Moreover, (3) JAGB stars can be easily and unambiguously identified on the basis of their near-infrared colors alone (without the need for spectroscopy or narrow-band photometry), thereby being readily distinguished from both the (bluer) O-rich AGB stars, as well as (even redder) “extreme” carbon stars.
Extending the JAGB method, Freedman & Madore (2020) used these stars to measure the distances to additional nearby galaxies within and beyond the Local Group, out to 27 Mpc. In a sample of 14 galaxies, they found excellent agreement of the JAGB distances with published TRGB distances to those same galaxies, where the combined scatter (including potential effects of metallicity and star formation history) amounts to only 4%. Simultaneously, Ripoche et al. (2020) investigated the Magellanic Clouds and the Milky Way. A number of additional extensive tests of this method have recently been carried out by Lee and collaborators Lee et al. (2021, 2022); Lee (2023); Lee et al. (2024a, b) as well as Parada et al. (2021) and Zgirski et al. (2021) in several nearby galaxies, confirming the excellent agreement with distances measured with the TRGB and Cepheid distance scales, and again indicating that any potentially confounding effects of metallicity variations and star formation history must be contained within that small scatter, and therefore must be relatively small, individually and collectively.
Two astrophysical effects are known to account for the small spread in the luminosity of carbon stars. Carbon stars are formed during the thermally pulsing phase for AGB stars when carbon is convectively dredged up to the surface of the star (Iben & Renzini (1983); Herwig (2013); Habing & Olofsson (2004); Marigo et al. (2017); Salaris et al. (2014); Pastorelli et al. (2020)). (a) The upper limit to the luminosity results from the fact that younger, more massive (hotter) AGB stars burn their carbon at the bottom of the convective envelope before it can reach the surface of the star Boothroyd et al. (1993). (b) The lower limit to the luminosity of carbon stars results from the fact that for the oldest, less massive AGB stars, there is no third (deeper) dredge-up phase. Thus, carbon stars emerge only in a well-defined intermediate mass range: where thermal pulses are effective dredging up carbon to the surface, but only when the carbon survives being burned at the lower levels before being mixed into the outer envelope.
The advantages of the JAGB method are many. (1) JAGB stars are the dominant population of the reddest stars found in galaxies. They are easily identified on the basis of their near-infrared colors. (2) The color-selected -band luminosity function is centrally peaked with a low dispersion. (3) JAGB stars are, in the mean, about one magnitude brighter than the brightest TRGB stars. (4) JAGB stars are found in all galaxies that have an intermediate-age population, and therefore the method can be applied to a wide range of galaxy types. (5) In the infrared, the total reddening along the line of sight (including Milky Way foreground extinction, host galaxy extinction, and any dust reddening generated by the carbon stars themselves) drops by about a factor of four in going from the optical () to the near infrared (). And, (6) unlike the use of period-luminosity relations for variable stars (e.g., Cepheids or Miras), which need to be monitored over multiple cycles, JAGB stars require only a single-epoch observation.
The disadvantages of the JAGB method are (1) Although minimized in the band, reddening variations within the host galaxy remain a source of uncertainty for the JAGB population. (2) Care must be taken to compare regions of comparable dust and metallicity levels in the calibrating and target galaxies. And (3) the comparative novelty of this method requires further study to quantify potential effects that might result, for example, from differences in star formation histories, or the existence of different population sub-types.
6.3 Cepheids
Cepheid variables have played a key role in the determination of extragalactic distances: over the last century, they became the de facto gold standard. For recent reviews of the Cepheid distance scale see Madore & Freedman (1991); Bono et al. (2010); Freedman & Madore (2010); Turner (2012); Freedman & Madore (2023a). Cepheid variability results from pulsation of their atmospheres, which is driven by a valve-like mechanism as helium cycles from a singly to a doubly ionized state, and the opacity increases as the star is compressed. The well-defined relationship between period, luminosity and color provides a standardizable candle of high precision. Including a metallicity term, the Leavitt law can be linearly parameterized as
where the Cepheid magnitude at a given wavelength is a function of the logarithm of the period (P) with coefficient , a color term with coefficient , and a term with coefficient , that allows for a metallicity dependence (where [O/H] represents the logarithmic oxygen-to-hydrogen ratio for HII regions in the vicinity of the Cepheids, relative to the solar value); and is the overall zero point.
A well-known characteristic of the PL relation is its decreasing scatter/width with increasing wavelength (e.g., Madore & Freedman, 1991). The simultaneous decrease of both reddening and metallicity effects, again with increasing wavelength, motivates the application of the Leavitt Law at longer (near- and mid-infrared) wavelengths (McGonegal et al., 1982; Madore & Freedman, 1991; Freedman et al., 1991; Macri et al., 2001; Freedman et al., 2008). (An exception is the 4.5m band in which the Cepheid flux is affected by the presence of a CO bandhead (Scowcroft et al., 2011)). Hence, the focus of recent applications to the Cepheid distance scale has been to discover these variables at optical wavelengths, and then following them up with available redder passbands (e.g., Freedman et al., 2001; Macri et al., 2001; Freedman et al., 2012; Riess et al., 2009, 2022).
If accurate colors are available, a reddening-free magnitude or Wesenheit function, , can be constructed, as first demonstrated by Madore (1976, 1982). is widely used in applications of the Cepheid distance scale (e.g., Freedman et al., 2001; Riess et al., 2016). For example, at optical wavelengths
where RV = AV /E(B-V) is the ratio of total-to-selective absorption. simultaneously corrects for total line-of-sight absorption, including both host-galaxy (internal) and Galactic (foreground) contributions.
The advantages of the Cepheid distance scale are many. (1) Cepheids are supergiant stars placing them among the brightest stellar indicators. (2) Cepheids are relatively young, high-mass stars, found in abundance in star-forming galaxies. (3) The Cepheid Leavitt relation, at the reddest wavelengths, has a small scatter (e.g., in the I or J band), where the dispersions amount to only 0.1 to 0.2 mag. (4) Cepheids are observationally straightforward to discover and uniquely identify. At optical wavelengths, they exhibit amplitudes up to 1 mag, and they display characteristic and identifiable (“sawtooth”) light curve shapes. And, (5) Cepheids can be re-observed over time (unlike supernovae), thereby allowing for the updating of their periods and phases, and adding additional wavelength coverage or achieving greater image resolution as new instrumentation becomes available, as, for example, in the case of HST and JWST.
There are also disadvantages of the Cepheid distance scale. (1) Cepheid variables are generally found in the relatively high-surface-brightness areas in the star-forming disks of late-type galaxies. In these regions crowding/blending effects increase the random and systematic noise seen in the individual Cepheid measurements and complicate sky subtraction. (2) Being relatively young stellar objects, Cepheids are located close to the regions of dust and gas out of which they formed. Hence, precise and accurate colors are required in order to correct the stars individually for the effects of reddening and extinction. (3) The effects of metallicity on the Cepheid Leavitt Law as a function of wavelength are still being actively debated in the literature (e.g. Ripepi et al., 2020; Madore & Freedman, 2024a). (4) Complexity. The many steps required to standardize Cepheid luminosities (measurement of light curves, noise threshold for inclusion in sample, completeness bias, adopted period range, artificial star corrections to correct for crowding/blending effects, extinction corrections, metallicity corrections) all involve choices. Cepheids are not simple standard candles.
7 New Distances from JWST Observations
As noted in §3, our observing program was configured to optimize simultaneous observations for all three of our methods555The JWST data for our program are available from the Barbara A. Mikulski Archive for Space Telescopes (MAST) Portal, Proposal ID 1995.. This goal led us to choose as our primary filter. First, JAGB stars have nearly constant luminosities in the () band. Second, observations provide an additional waveband beyond the HST and filters, to improve the extinction correction for Cepheids. Third, extinction effects are small at 1.15 m, which is an advantage for the TRGB method relative to the band (where extinction effects are about a factor of two larger). We note that the disadvantage of the -band for the TRGB method is that there is a steeper color dependence than for the band where the TRGB is relatively flat (e.g., Lee et al., 1993). Nonetheless, the TRGB is still standardizable in the near-infrared (e.g., Dalcanton et al., 2012; Madore et al., 2018; Hoyt et al., 2018; Durbin et al., 2020; Hoyt et al., 2024).
As also noted in §3, our original choice for our second filter was motivated to allow us to provide constraints on the metallicity sensitivity of the Leavitt law. At 4.5 m there is a metallicity dependence of the Leavitt law as a result of a CO-band head (Scowcroft et al., 2011). The filter can also provide a broad color baseline for the JAGB method. However, the relatively low resolution of the data prompted us to switch to the filter, which provides both higher resolution and higher sensitivity.
7.1 TRGB Measurements
An important aspect of an accurate measurement of an old TRGB population is the spatial selection – finding an optimal balance between avoiding regions of the disk where younger populations reside, and not being so far out in the halo so that the number density of stars is too small. As described in H24, we used the blue, unresolved light to trace regions dominated by young disk populations, masking regions bright in blue flux.
In Figure 4, we show examples of versus ( - ) color-magnitude diagrams for three of our program galaxies, NGC 4258, NGC 4424 and NGC 4039 from H24. The position of the TRGB is shown, and is (both qualitatively and quantitatively) easily identified. As can be seen, the TRGB feature in the infrared exhibits a slope with photometric color. This dependence is due to a combination of metallicity and, to a lesser extent, to age effects. In the band, the metallicity dependence is about five times larger than that due to age (e.g., McQuinn et al., 2019). To measure the TRGB in each SN Ia host, the gradient in the magnitude direction was computed from a smoothed Hess diagram of the selected RGB stars, and the centroid of the dominant contour defined the TRGB in each field.
Figure 5 illustrates a smoothed Hess diagram of RGB stars in the galaxy NGC 4424. For each galaxy, the magnitude-direction gradient contained a maximal set of contours that indicated the location of the TRGB. A threshold between the 90th and 99th percentile of the 2-D gradient was then used to define the TRGB contour. An ellipse was then fit to this TRGB contour and the centroid of that ellipse was adopted as the best-fit magnitude and color of the TRGB. To compute the measurement uncertainty, the length of the axis perpendicular to the sloped TRGB was scaled by the contrast of the TRGB ellipse, which was computed as the ratio of the number of stars contained on the faint end of the semi-major axis chord to those above it. The mean TRGB color and magnitude estimated for each SN Host and for three fields in NGC 4258 were then combined with an estimate of the TRGB color slope (see H24 for details) to determine distance moduli. With a mean TRGB color and magnitude estimated for each SN Ia host, as well as for each of three fields in NGC 4258, distance moduli were computed relative to NGC 4258.
7.2 JAGB Measurements
In brief, as described in detail in L24, the JAGB stars in our target galaxies were selected based on their position in (initially blinded) near-infrared () or () JWST NIRCam color-magnitude diagrams. The magnitudes were binned and the luminosity functions were smoothed using a Gaussian-windowed, Locally Weighted Scatterplot Smoothing (GLOESS) algorithm (e.g., Persson et al., 2004). The JAGB magnitudes for each galaxy were determined from the mode of the smoothed luminosity function.
The JAGB method is best applied in the outer disks and halos of galaxies where there are sufficient numbers of carbon stars to provide a statistically meaningful measurement, but not too far into the disk where systematic effects from crowding, blending, and reddening become an issue, and the luminosity function becomes asymmetric. As demonstrated in L24, the JAGB luminosity exhibits a well-defined mode and a distinct Gaussian form in the lower-reddening, less crowded regions of the galaxies. (The mode is also robust against asymmetries in the underlying distribution, as well as robust against outliers in the tails and/or windowing or clipping.) The value of the mode is generally found to be brightest in the inner, high-surface brightness regions of a galaxy where the crowding and reddening effects are greatest. For a clean measurement of the JAGB luminosity function, these high surface-brightness regions need to be avoided.
The radial cuts for the JAGB luminosity functions were set by seeking convergence in the radial distribution: that is, the radial distance, r, within the disk where the derivative of the magnitude of the JAGB (d/dr) stabilized and leveled off to zero (L24). In three cases (NGC 3972, NGC 4424 and NGC 4038), no clean convergence was found, and these galaxies were not included in further analysis.
7.3 Tests for Systematics in Distances Measured Using the JAGB Method
To test how our choice of JAGB statistic affected the final measured distances, we explored how the smoothing parameter affected the final measured mode, since the mode measured from an increasingly smoothed luminosity function will eventually converge to the mean. We varied the smoothing scale of the JAGB luminosity function using smoothing scales of (0.15, 0.20, 0.25, 0.30, 0.35, 0.40) mag, and then re-measured the mode with the given smoothing scale. We defined the statistical error due to the choice of smoothing parameter as the maximum difference between the fiducial mode (measured with a smoothing parameter of mag for all galaxies) and all of the measured modes. The total smoothing parameter error was then defined to be the smoothing parameter error from the SN Ia host galaxy and the zeropoint from NGC 4258 added in quadrature. This error was larger than 0.05 mag for all galaxies. In Lee+ (2024, submitted), we demonstrated that using the mean/median instead of the mode as the chosen JAGB statistic resulted in distance moduli that were measured to be 0.03 mag brighter on average. Thus, this systematic offset was fully encapsulated within the minimum adopted smoothing parameter uncertainty of 0.05 mag. Differences between the mean and the modal values of the JAGB populations were also explored in the study of Madore et al. (2022); see their Figure 1.
The effect of a variety of other terms potentially affecting the JAGB distance determination method have been discussed previously in detail in Freedman & Madore (2020). These include the effects of star formation history, C-to-M AGB ratio variations, metallicity variations, mass loss, JAGB star variability, foreground/background contamination and lastly, the effect of moving the color-selection window.
For our current study, we adopted a statistical error due to the fluctuations about the final converged JAGB magnitude past the adopted outer disk radial cut. This uncertainty was derived from the dispersion about all measured outside of the radial cut, divided by the square root of the number of bins. The presence of a spiral arm in the convergence plots sometimes caused additional noise. As a result, converged at a radial distance with too few JAGB stars or dm/dr not equal to zero in four galaxies: M101, NGC 2442, NGC 4258, and NGC 4639. We masked these spiral arms, and then re-calculated the convergence plots, where then successfully converged. We emphasize we performed this procedure during the blinded stage of our analysis. We only masked the spiral arms so that successfully converged, independent of knowledge of the final measured distance. However, leaving the spiral arms unmasked while using the newly adopted radial cuts yielded almost a negligible change in of 0.3% (larger).
7.4 Cepheid Measurements
At present, measurement of the Cepheid distances cannot be obtained from our JWST data alone: to correct for extinction necessitates measurements at multiple wavelengths. For this purpose we used existing HST/ archival data at optical ( and ) wavelengths obtained by the SHoES team. We independently reduced all of the archival data for our target galaxies.
Initial Cepheid candidates were obtained directly from Riess et al. (2016, 2022). Periods and mean magnitudes for these candidates were obtained using template light curves from O24. We eliminated candidates with poor light curve fits or extremely low amplitudes (i.e., 0.2 mag in the or band: the filter used was that upon which the published periods were based), as well as candidates with periods in excess of 100 days. Examples of PL relations at , , and NIRCam are shown in Figure 6 for the galaxies NGC 5643 and NGC 4639. In order to correct the magnitudes for crowding bias, we carefully injected 1,000 artificial stars, one at a time, within a 3 arcsec radius of each Cepheid. Following the results of the artificial star tests, we eliminated Cepheids that were close to the edge of the NIRCam frames, as they had lower quality photometry due to non-overlapping dithers. We fit period-luminosity (PL) relations using fixed slopes from Monson et al. (2012) in the NIRCam , , , and bands. Finally, we excluded obvious outliers based on a 4-sigma clip.
For our primary analysis, we calculated PL relation intercepts (and their uncertainties) through resampling, including redrawing from the distribution of the artificial star corrections for each Cepheid. The PL relation intercepts and uncertainties were taken to be the mean and standard deviation of the simulated distribution. We fit a Cardelli et al. (1989) reddening law with Rv = 3.1 to the set of resampled PL relations, extrapolating to infinite wavelength, and adopted the mean and standard deviation of the distribution of reddening law fits as the absolute distance modulus and statistical uncertainty. In addition, we carried out two additional cross-checks on the distances. We also directly measured PL relations and reddening law fits without weighting or resampling, adopting the mean artificial star corrections for each Cepheid and for each bandpass. In addition, we constructed Wesenheit PL relations (Madore, 1976) using the magnitudes and () colors to measure unweighted reddening-corrected distances. The median offset of the absolute distance moduli compared to the primary analysis for the 10 SN Ia host galaxies was 0.01 mag for both comparisons.
Following on from our blind analysis, we also undertook a broader analysis of the fitting of the Cepheid PL relations. In this case, we employed a Sobel edge-detector to identify the upper and lower 2- edges of the instability strip that define the boundaries of the PL relations. This was done by marginalizing the PL relations using previously determined fixed slopes. The intrinsic widths of the PL relations as a function of wavelength were obtained from well-observed Cepheids in the LMC. This method was found to be quite insensitive to choice of period cuts and incomplete sampling within the instability strip. It also provided a means of identifying outliers. It yielded significantly larger distance moduli, primarily a function of the fact that within our calibrating galaxy, NGC 4258, there are few faint stars with short periods, as well as few bright Cepheids with longer periods, thereby leading to a significant flattening of the PL relations due to the underpopulated extremes of the PL relations.
One can, of course, make a series of choices in how to define the Cepheid sample selection, whether to impose period cuts, how to measure the extinction, and what, if any, metallicity correction is required. As our independent analyses have revealed, one can make reasonable (defensible) choices, but obtain significantly different results. Moreover, it is non-trivial to quantitatively assess the impact of these choices on the error budget. Unfortunately, these subtle data-reduction choices can determine whether or not there is an tension. Our view is that these discrepant values provide valuable insight into the current uncertainties. We leave to future studies (with improved colors, and smaller scatter in the PL relations) a resolution of these discrepancies. Because running our PL fits with various period cuts leads to significantly different distance moduli, at this juncture, we conservatively adopt an additional uncertainty of 0.07 mag for our Cepheid distances.
8 Comparison of the Cepheid, TRGB and JAGB Distances
In Table 2, we list the supernovae, host galaxy name, individual distances plus their uncertainties for each of the methods. Detailed descriptions of the measurement of these distances are provided in O24, H24 and L24. Here we compare the three measurements. We note that for the galaxy NGC 5643, there are two SNe Ia that have been observed, SN 2013aa and SN 2017cbv. For the Cepheid and TRGB calibrations, there are 11 SN Ia calibrators that have been observed as part of Pantheon+, and 10 for the CSP. (The CSP program ended in 2015 before SN 2017cbv was discovered.) For the JAGB calibration there are ten SN Ia calibrators, but three of these are of significantly lower weight. As described in L24, for three of the galaxies in our sample (NGC 3972, NGC 4038, and NGC 4424), no convergence in the radial measurement of the luminosity function was found.
SN | Galaxy | ||||||
2011fe | M101 | 29.18 | 0.04 | 29.14 | 0.08 | 29.22 | 0.04 |
2012fr | N1365 | 31.33 | 0.07 | 31.26 | 0.1 | 31.39 | 0.04 |
2015F | N2442 | 31.61 | 0.09 | 31.47 | 0.09 | 31.61 | 0.04 |
2011by | N3972 | 31.74 | 0.07 | 31.67 | 0.10 | … | … |
2007sr | N4038 | 31.61 | 0.08 | 31.70 | 0.12 | … | … |
2012cg | N4424 | 30.93 | 0.05 | 30.91 | 0.22 | … | … |
1981B | N4536 | 30.94 | 0.06 | 30.95 | 0.12 | 30.98 | 0.03 |
1990N | N4639 | 31.75 | 0.07 | 31.8 | 0.12 | 31.74 | 0.04 |
2013aa | N5643 | 30.61 | 0.07 | 30.51 | 0.08 | 30.59 | 0.04 |
2017cbv | N5643 | 30.61 | 0.07 | 30.51 | 0.08 | 30.59 | 0.04 |
2013dy | N7250 | 31.62 | 0.04 | 31.41 | 0.12 | 31.6 | 0.08 |
An immediate result from our comparisons of the three methods is that the TRGB and JAGB distances are in superb accord, whereas the measured Cepheid distances are offset (closer) than the other two. The upper panel of Figure 7 compares JAGB and TRGB distance moduli in our sample. The weighted (unweighted) mean difference between the JAGB minus TRGB distance moduli is 0.017 0.031 (0.013 0.013) mag or 1%. The scatter about the unit slope is 0.031 mag (1.4%). Immediately below, the residuals from the unit slope line are shown.
In the middle panel of Figure 7 we compare the Cepheid and TRGB distance moduli in our sample. In this case, the agreement is not as good, but is still reasonable: the scatter about the unit slope is 0.083 mag (3.9%). The weighted (and unweighted) mean difference between the Cepheid minus TRGB distance moduli is -0.054 0.028 (-0.049 0.028) mag, or a difference of about 2.5%.
Finally, the lower panel of Figure 7 compares the individual JAGB and Cepheid distance moduli in our sample. The scatter about the unit slope is 0.070 mag (3.3%). The weighted (unweighted) mean difference between the JAGB minus Cepheid distance moduli is 0.086 0.028 (0.083 0.031) mag or 4%.
We note that our Cepheid distances are in good agreement with those measured by R22. In Figure 8, we show a comparison of Cepheid distances measured in this study with those of R22. In this case, the weighted mean offset is -0.022 mag, amounting to a difference in distance (and ultimately ) of only 1%. The difference is in the sense of the new JWST distances being smaller.
The agreement for two of the (independent) methods (the JAGB and TRGB distances) is encouraging. The implication of these comparisons is that the TRGB and JAGB calibrations then lead to a lower value of than for the Cepheids; i.e., there is a local distance scale tension, independent of the tension in cosmology. These methods are all being applied to the same galaxies, for which there is a single distance. Understanding the reasons for these offsets is essential before we will be able to confirm if the Hubble tension is real. Importantly, we note that these differences arise in the distances themselves, before any application to the SNe Ia distance scale or determination of .
In concluding the discussion in this section, we note that there are several improvements to the analysis of Cepheid distances obtained as part of this study over those measured with HST alone, but in addition, there is a disadvantage. The advantages include the fact that (1) the new JWST near-infrared data have higher S/N (by more than an order of magnitude) than the previous HST near-infrared data. (2) The resolution is increased by a factor of four. (3) The re-analysis of the optical HST data has the benefit of the input source catalog coming from JWST. Hence, more stars contaminating the Cepheids can be removed in these new reductions. (4) As described in more detail in O24, and illustrated in Figure 9, three bands (F555W, F814W and F115W) provide a well-delineated extinction curve from which a correction for dust and a (total) reddening-corrected distance modulus can be obtained (e.g., Freedman, 1988; Freedman et al., 1991; Monson et al., 2012; Scowcroft et al., 2013). The dust corrections are consistent with what is expected for the foreground reddenings to these target galaxies, plus an additional contribution from dust internal to the target galaxy. However, a major disadvantage of the current JWST data is the smaller number of Cepheids observed in the calibrating galaxy, NGC 4258. The sampling within the Leavitt relation is such that a classical (least-squares) fit to the PL is very sensitive to the choice of short-period cut to the data. The uncertainty in the Cepheid distances for this calibration are thus larger than those in R22.
In §9 we provide a determination of based on each of the three distance scales individually, as well as a combined value based on the three methods.
9 Calibration of SNe Ia and the Hubble Constant
None of the three methods considered here (Cepheids, TRGB or JAGB stars) are bright enough to determine distances out into the smooth cosmic Hubble flow at a level of 1% accuracy in . Nearby galaxies and clusters induce motions or peculiar velocities, scattering above and below the Hubble expansion velocities, adding noise and potential bias to the determination of . In the past couple of decades, SNe Ia have surfaced as the preferred secondary distance indicator given their high intrinsic brightness and their small observed dispersion in the Hubble diagram (0.1-0.15 mag, e.g., (Burns et al., 2018; Scolnic et al., 2022; Brout et al., 2022)). We apply our independent Cepheid, TRGB and JAGB star distances to two samples of distant supernovae: 1) the ( CSP) (Uddin et al., 2024) and 2) the Pantheon+ sample (Scolnic et al., 2022), as described below.
9.1 The Carnegie Supernova Program (CSP)
The CSP was initiated 20 years ago as a program to provide multi-wavelength follow-up observations of previously discovered supernovae. The data were intended for applications to cosmology, as well as for studying the physical properties of the supernovae themselves666The CSP data are available at http://csp.obs.carnegiescience.edu/data.. The aim was to obtain homogeneous, intensive, high-cadence, multi- observations (Contreras et al., 2010). Careful attention was given to photometric precision and systematics: the program utilized a fixed set of instruments, photometric standard stars, and instrumental reduction procedures, catching most of the supernovae well before maximum, and with high signal to noise, avoiding many of the challenges otherwise faced in minimizing systematic differences between multiple data sets/instruments/etc. (Krisciunas et al., 2017). Optical spectra were also obtained with high cadence (Folatelli et al., 2013; Morrell et al., 2024). The bulk of the observations were carried out at Las Campanas Observatory using the 1-m Swope and 2.5-m du Pont telescopes. The first part of the CSP ( CSP-I) was carried out from 2004-2009, with a second phase ( CSP-II) from 2011-2015, optimized for the near-infrared (Phillips et al., 2019; Hsiao et al., 2019). Light curves were generated using the analysis package SNooPy (Burns et al., 2018). This program determines the peak magnitude in the light curves for each of the filters, the times of those maxima, (B-V) colors, and a color-stretch parameter sBV, as described by (Burns et al., 2014). The color-stretch parameter incorporates color in the relation between the luminosity and decline rate of a SN Ia, allowing also for the redder colors of fast decliners. Finally it computes all of the covariances amongst the parameters.
Previous applications of the CSP-I survey to cosmology include Burns et al. (2018), who used the CSP-I SNe Ia sample, calibrated using the Cepheid distances of Riess et al. (2016), to obtain a value of = 73.2 2.3 km s-1 Mpc-1 (for H-band data); and a value of = 72.7 2.1 km s-1 Mpc-1 (for B-band data). Calibrating the CSP-I sample with the TRGB (Freedman et al., 2019), updated in (Freedman et al., 2020; Freedman, 2021) found a lower value of = 69.8 0.6 (stat) 1.6 (sys) km s-1 Mpc-1. Uddin et al. (2024) have updated the CSP analysis, including the more recent CSP-II SNe Ia data, which, once included with CSP-I, triples the sample size over CSP-I. They undertook a calibration based on Cepheids, the TRGB and Surface Brightness Fluctuations, the latter a secondary distance indicator calibrated first by Cepheids. Using B-band light-curve fits, they find = 73.38 0.73 km s-1 Mpc-1 based on the Cepheid calibration of Riess et al. (2022). For the TRGB calibration, they find = 69.88 0.76 km s-1 Mpc-1, based on the TRGB distances of Freedman et al. (2019). Both of these results are in excellent agreement with the original studies.
9.2 JWST CSP calibration
The SNe Ia data used in this work are those described in the recent analysis by Uddin et al. (2024). There are over 300 SNe Ia in this sample including both CSP-I and CSP-II at wavelengths, and more than 200 SNe Ia observed at . In analyzing these data, we follow the methodology described previously by Burns et al. (2018); Freedman et al. (2019); Uddin et al. (2024). As in Burns et al., the corrected magnitudes for an individual filter (e.g., the B filter below) are given by:
where in this case = , the apparent peak (K-corrected) magnitude, is the linear coefficient and is a quadratic coefficient in (); is the slope of the color correction; is the apparent peak magnitude at , K-corrected; and is the slope of the correlation between peak luminosity and host stellar mass . Host stellar masses are derived as described in Uddin et al. (2024). Following Uddin et al., the sample is split at the median mass so that equal weights are given above and below the median mass (although the results are not significantly affected by the choice of split point).
The apparent magnitudes at maximum are computed by fitting the light-curves with SNooPy, providing the time of maximum, the light-curve shape , and the magnitude at maximum for each filter. These quantities are then provided as inputs to a Markov Chain Monte Carlo (MCMC) sampler that simultaneously solves for all the correction factors: , , , and (for full details, see Burns et al., 2018; Uddin et al., 2024). The MCMC sampler then provides the corrected magnitudes, as well as a full covariance matrix, which is used when determining and its error.
The distance modulus is defined as in (Burns et al., 2018) for a flat cosmology:
where is the deceleration parameter.
As described in Uddin et al. (2024), three error terms are included in the analysis. The first, is the sum of the individual errors for the observed quantities, in addition to the covariance between the peak magnitude, and both the color and the color-stretch parameter. The second, is the intrinsic random scatter that allows for variations beyond either or those due to peculiar velocities, . The third term is the error due to the uncertainty in the distance measurements resulting from galaxy peculiar velocities, and scales with redshift. As in Burns et al. (2018), this term is given by where Vpec is a free parameter that represents the average peculiar velocity of the SN Ia sample.
There are some differences in the CSP-I and CSP-II samples previously noted by Uddin et al. (2024). The scatter in the CSP-II data is larger than for CSP-I, which as noted by Uddin et al., results from a doubling of the number of low-redshift SNe Ia in the combined CSP sample. The average peculiar velocity (vpec) for the CSP-II sample is also larger than found for CSP-I, a consequence of increasing the number of low-redshift SNe Ia in the combined sample. The CSP-II sample doubled the total number of low-redshift SNe Ia with z (with 0.01 (14 SNe Ia in CSP-I, and 13 SNe Ia in CSP-II). To avoid issues with the largest peculiar velocities for nearby objects, in this paper, we limit the redshift to z 0.01 (e.g., see Brout et al., 2022). The parameter is also steeper in the combined sample than the CSP-I sample. One possibility is that the combined sample could contain more star-forming galaxies with larger amounts of dust. These issues highlight the fact that in the future, in an era where a value of at a level of 1% accuracy is the goal, large and homogeneous samples of SNe Ia become imperative.
9.3 The Hubble Diagram
In Figure 10 we show the Hubble diagram for the CSP sample of 287 SNe Ia (blue filled circles ( CSP-I) and dark cyan filled circles ( CSP-II) where z0.01 and excluding super-Chandra (IA-SC) and IaX subtypes. In addition, two additional 3- outliers (SN2014D and SN2013hh) were excluded from the analysis. The calibrating galaxies provide distances only; their velocities are not used in the calibration of . Residuals from the fit in distance modulus are shown in the lower panel.
9.4 Markov Chain Monte Carlo Analysis
We explore the use of two Python packages, emcee (Foreman-Mackey et al., 2013), and pymc3 (Oriol et al., 2023) and find that they give consistent results. Initially, using emcee, a set of broad, uniform priors were adopted for 8 parameters (P0, P1, P2, , , Vpec, , and ) and 100 walkers were used. The burn-in time was set to be 5 times the autocorrelation time, , generally 500 steps, and the acceptance fraction was generally 0.45-0.49, within the range of 0.2-0.5 recommended by Gelman et al. (1996). The number of chains after burn-in was taken to be greater than 100, and convergence was visually checked in walker trajectory plots. The adopted parameter fits are given by the marginalized distributions and uncertainties quoted are 16th, 50th and 84th percentiles. In the case of the second analysis using pymc, 30,000 steps were implemented, and the burn-in time was set to 3,000. pymc uses “forward sampling” using the derivatives of the probability hyper-surface. Consistent results were obtained in both cases.
We note that the sample of ten galaxies for which we have JWST data/distances for is currently small, resulting in a significant statistical uncertainty relative to longer-running HST-based studies (e.g., F19, R22). The scatter in the peak SN Ia magnitudes for this nearby sample is smaller than for the distant CSP sample. Allowing only for a single parameter to account for the scatter, , leads to an anomalously small dispersion in the P0 posterior, and hence, the posterior. We thus included an additional 9th parameter (), allowing for this difference in the scatter between the calibrating and distant sample, and thereby allowing for a larger uncertainty in . The scatter in P0 then increased by a factor of five, and the resulting value of increased by 1%, with a two-fold increase in its uncertainty. At present, it is not clear if the larger scatter for the more distant sample is a result of larger (unaccounted-for) errors (for example, errors in k-corrections, which are currently not included). Alternatively, the nearby sample may not be large enough to sample the true scatter. This situation will become clarified as further JWST data become available.
In Table 3 we list the parameter fits individually for the pymc runs for the Cepheid, TRGB and JAGB distances. We show an example corner plot from pymc for the TRGB calibrators in Figure 13. From our adopted pymc analysis, based on the -band CSP SN Ia data, the peak of the marginalized distributions for the three individual distance indicators are: = 72.05 1.86 for Cepheids, = 69.85 1.75 for the TRGB, and = 67.96 1.57 km s-1 Mpc-1 for the JAGB (statistical errors). These values are based on 10 SN Ia calibrators for the TRGB and Cepheids, and 7 calibrators for the JAGB method. The estimated statistical uncertainty calculated for the JAGB method (and listed in Table 3) is lower than that for the Cepheids and the TRGB simply as a consequence of the fact that the blinded sample of seven calibrating galaxies does not include three additional supernovae, SN 2011by, SN 2007sr, and SN 2012cg in the galaxies NGC 3972, NGC 4038 and NGC 4424, respectively. As described in §7.2, the radial measurements of the JAGB magnitude did not converge in these galaxies. In order not to give undue weight to the JAGB method, we increased its formal uncertainty (1.57) by the ratio of the average values (based on the TRGB and Cepheids) to that of the lower value for the JAGB: i.e., 0.157 0.13/0.11 = 1.85. A Bayesian analysis combining the three methods yields a value of = 69.96 1.05 km s-1 Mpc-1 (statistical error). We discuss the systematic uncertainty in §11.
Parameter | TRGB | JAGB | Cepheids |
P0 (mag) | -19.20 0.05 | -19.26 0.05 | -19.13 0.05 |
P1 (mag) | -0.90 0.10 | -0.90 0.10 | -0.90 0.10 |
P2 (mag) | -0.28 0.29 | -0.29 0.29 | -0.29 0.28 |
2.91 0.09 | 2.92 0.09 | 2.92 0.09 | |
(mag/dex) | 0.00 0.01 | 0.00 0.01 | 0.00 0.01 |
(mag) | 0.14 0.05 | 0.11 0.05 | 0.12 0.06 |
(mag) | 0.19 0.01 | 0.19 0.01 | 0.19 0.01 |
Vpec (km s-1) | 174 99 | 175 99 | 175 99 |
H0 | 69.85 1.75 | 67.96 1.57aaAs described in the text, when combining the methods for the final analysis, the uncertainty for the JAGB method was increased by a factor of 1.18 (given by the ratio of the average values for TRGB and Cepheids to that of the lower value for the JAGB), so that the JAGB method does not unduly contribute more weight than the other two methods. | 72.05 1.86 |
As previously noted, the total CSP sample includes a significant fraction of low-z SNe Ia which, if included in the MCMC analysis, would require a correction for the local density field. For example, based on predictions from the 2M++ density field of Carrick et al. (2015), Uddin et al. (2024) corrected the CSP peculiar velocity measurements (for which the average peculiar velocity is 440 km s-1). They found a mean sample velocity correction for the CSP survey of 90 km s-1, resulting in a net increase in of 0.55 km s-1 Mpc-1 in the B band. However, as shown in Brout et al. (2022), not including supernovae at z 0.01 avoids the bias due to nearby peculiar velocities. Also as previously noted, our preferred value of is based on a sample of 287 SNe Ia with z0.01, eliminating also super-Chandra (IA-SC) and IaX subtypes, thus requiring no additional correction to .
In Table 4 we summarize the values of obtained using the three methods applied in this paper, based on the CSP SN Ia data. We show also values that are based on the Pantheon+ SN Ia data (see discussion in §9.5 below), for which there is good agreement. We present the results from different MCMC analyses (pymc and emcee), show the results adopting different redshift cuts for the CSP SNe Ia sample, adopting different nearby calibrators, calibrating the -band data for the CSP SNe Ia, and, for the Cepheids, with and without metallicity corrections. To within the uncertainties, the various cuts do not lead to significant variations in . The largest differences remain between the Cepheid compared to the TRGB and JAGB calibrations. We note that the total CSP sample of SNe Ia with -band data is not as large as that for the band (213 compared to 322), and the values of obtained based on the -band are larger than for the band. Uddin et al. (2024) searched, but found no evidence, for potential calibration errors. There are, however, hints of the existence of a brighter sub-population of SNe Ia in the near-infrared, which warrants further investigation. Finally, we list the combined, adopted result. The uncertainties in these values are discussed in §11.1.
Our adopted value of , based on the more extensive -band data is listed at the end of the table. We show the relative probability density functions (PDFs) for the three methods and the combined, adopted value in Figure 11. A Frequentist cumulative sum of the PDFs yields = 69.82 1.69 (stat) 1.12 (sys). A Bayesian analysis based on the product of the PDFs, adopting a flat prior, yields = 69.96 1.05 (stat) 1.12 (sys). A discussion of the uncertainties adopted follows in Table §5. We consider below the combined value of based on the total errors, but our preference for our adopted value is to keep the statistical and systematic errors separate, and easily identified.
In Figure 12 we plot the relative probability density functions for the three methods taking into account both the statistical and systematic uncertainties. In this case, the width of each Gaussian is given by . Assuming a flat prior, the Bayesian product of the PDFs yields = 69.59 1.58 (2.3%) km s-1 Mpc-1, based on all three methods (black curve). The JAGB and TRGB methods alone yield = 69.03 1.75 (2.5%) km s-1 Mpc-1 (blue-gray curve). We note that the latter is based solely on JWST data: that is, the calibration through NGC 4258 to the target galaxies is carried out with the same NIRCam detector. The Cepheid analysis relies on both HST and JWST data.
aafootnotemark: | Erroraafootnotemark: | Descriptionb,cb,cfootnotemark: | MCMC code |
TRGB | |||
68.63 | +1.38 -1.32 | TRGB + CSP B band ; 9 SN calibratorsbbfootnotemark: , z0.01 | pymc |
68.62 | +1.41 -1.33 | TRGB + CSP B band ; 9 SN calibrators, z0.0065ccfootnotemark: | pymc |
68.72 | +1.38 -1.30 | TRGB + CSP B band ; 9 SN calibrators, z0ccfootnotemark: | pymc |
69.85 | +1.80 -1.70 | TRGB + CSP B band ; 10 SN calibrators, z0.01 | pymc |
68.03 | +1.48 -1.34 | TRGB + CSP B band ; 7 SN calibrators, z0.01 | pymc |
69.74 | +1.77 -1.71 | TRGB + CSP B band ; 10 SN calibrators, each with equal weight, z0.01 | pymc |
68.46 | +1.45 -1.34 | TRGB + CSP B band ; 9 SN calibrators, z0.01 | emcee |
69.64 | +1.73 -1.66 | TRGB + CSP B band ; 10 SN calibrators, z0.01 | emcee |
69.88 | +1.78 -1.74 | TRGB + CSP B band ; 10 SN calibrators, z0ccfootnotemark: | emcee |
JAGB | |||
68.76 | +1.86 -1.76 | JAGB + CSP B band ; 9 SN calibratorsddfootnotemark: , z0.01 | pymc |
68.92 | +1.64 -1.58 | JAGB + CSP B band ; 10 calibrators, z0.01 | pymc |
67.96 | +1.60 -1.53 | JAGB + CSP B band ; 7 SN calibrators, z0.01 | pymc |
Cepheids (z0.01; CSP B band ) | |||
70.77 | 1.55 -1.55 | Cepheids (with metallicity correction) ; 9 SN calibrators | pymc |
70.76 | 1.55 -1.56 | Cepheids (with metallicity correction) ; 9 SN calibrators, z0.0065ccfootnotemark: | pymc |
70.82 | 1.57 -1.56 | Cepheids (with metallicity correction) ; 9 SN calibrators;, z0ccfootnotemark: | pymc |
72.05 | 1.85 -1.86 | Cepheids (with metallicity correction) ; 10 SN calibrators | pymc |
72.56 | 2.00 -2.03 | Cepheids (without metallicity correction) ; 10 SN calibrators | pymc |
71.35 | +1.50 -1.58 | Cepheids (without metallicity correction) ; 9 SN calibrators | pymc |
Combined Results (z0.01, CSP B-band) | |||
69.82 | 1.69 1.12 | Combined (Cepheids, TRGB, JAGB) ; 10, 10, 8 SN calibrators; Frequentist | pymc |
69.96 | 1.05 1.12 | Combined (Cepheids, TRGB, JAGB) ; 10, 10, 7 SN calibrators; Bayesian | pymc |
69.03 | 1.75 | Combined (TRGB, JAGB alone) ; 11, 7 SN calibrators; Bayesian | pymc |
Pantheon SNe Ia sample | |||
69.94 | … | TRGB + Pantheon+ B band ; 11 SN calibrators | SHoES eefootnotemark: |
68.56 | … | JAGB + Pantheon+ B band ; 8 SN calibrators | SHoES eefootnotemark: |
71.95 | … | Cepheids + Pantheon+ B band ; 11 SN calibrators | SHoES eefootnotemark: |
H-band Results | |||
74.71 | +1.34 -1.35 | Cepheids CSP H band; 10 SN calibrators, z 0.01 | |
71.59 | +1.85 -1.76 | TRGB + CSP H band ; 10 SN calibrators, z 0.01 | |
70.96 | +1.91 -1.81 | TRGB + CSP H band ; 9 SN calibrators, z 0.01 | |
69.34 | +1.71 -1.62 | JAGB + CSP H band ; 7 SN calibrators, z 0.01 | |
Adopted Result | |||
69.96 1.05 (stat) 1.12 (sys) |
In summary, the final value that we adopt for this paper is based on all three calibrators, tying in to the -band CSP SN Ia sample. As noted, there is a systematic offset between the Cepheid distance scale and the JAGB/TRGB distance scale. We have listed separately the values obtained for the combination of TRGB and JAGB calibrations, and that from Cepheids. It will be important for future studies to resolve this local distance-scale discrepancy.
9.5 Comparison with Pantheon+
In the Pantheon+ analysis, Scolnic et al. (2022) collected and cross-calibrated the data for 1550 individual SNe Ia, superseding earlier Pantheon (Scolnic et al., 2018) and Joint Light-Curve (Betoule et al., 2014) analyses. The total sample includes SNe Ia in the redshift range 0 z 2.3. The analysis aims to standardize the B-band photometry from 18 individual surveys obtained with a wide variety of telescopes and instruments777 The Pantheon+ catalog is available at https://github.com/PantheonPlusSH0ES/DataRelease.. In the SHoES determination of , (e.g., Riess et al., 2009, 2022), a simultaneous fit is undertaken for the Cepheid and SNe Ia data, minimizing a statistic, and most recently providing the covariances. R22 also utilize emcee as a check of their methodology, finding almost exact agreement. The SHoES Cepheid calibration of the Pantheon+ SNe Ia sample from Scolnic et al. (2022) results in a value of = 73.04 1.04 km s-1 Mpc-1 for the 277 SNe Ia with 0.023 z 0.15.
In Figure 14 we compare the absolute magnitude SN Ia peak brightness, MB, for the set of SNe Ia contained in the Pantheon+ sample for which there are R22 Cepheid distances, and that now have distances also measured based on the three methods from this study. There are 10 galaxies for which we have measured TRGB and Cepheid distances (hosting 11 SNe Ia), and 7 galaxies with well-measured JAGB distances (hosting 8 SNe Ia). The weighted mean magnitudes and error on the mean for the three samples are = -19.286 0.029 mag (Cepheids), -19.345 0.026 mag (TRGB), and -19.384 0.023 mag (JAGB), respectively. These values can be compared to the overlapping sample of R22 SN Ia discussed above where, = -19.315 0.032 mag.
For comparison purposes, to be completely consistent with the SN Ia analyses undertaken by Scolnic et al. (2022) and R22, we do not re-analyze the Pantheon+ data. We determine by scaling the absolute MB values relative to those measured by SHoES, allowing for the fact that our calibration is based on NGC 4258 alone, and not the baseline calibration of the LMC, Milky Way and NGC 4258. The R22 value of based on NGC 4258 alone is 72.51 1.54 km s-1 Mpc-1 (which can be compared to their baseline value of 73.04 1.01 km s-1 Mpc-1). In the case of the Pantheon+ data we find = 71.95 Cepheids, = 69.94 for the TRGB, and = 68.56 km s-1 Mpc-1 for the JAGB. These values are in excellent agreement with those determined based on the CSP SNe Ia calibration.
10 Uncertainties in the MB Calibration for SNe Ia
A current and major uncertainty in the SN Ia distance scale is the relatively small number of SNe Ia calibrators in the local universe that are accessible to HST where Cepheid, TRGB or JAGB distances can be measured. At present, there are 42 SNe Ia located in 37 galaxies where Cepheids have been discovered by the SHoES team. In Figure 15 we plot the absolute peak B magnitudes, MB, for this sample of 42 SNe Ia as a function of distance modulus, all based on the Cepheid distance scale and SN Ia magnitudes from R22. All of the data points (blue and orange) are from the R22 data set. The orange points are the subset of SNe Ia observed as part of this study. The JWST sample selection was based on proximity, to minimize any possible systematics due to crowding/blending effects.
A 3 trend can be seen in this figure, in the sense that the supernova absolute magnitudes become fainter with increasing distance. In addition, for the nearest third of the SN Ia sample, the average MB magnitude is -0.07 mag brighter than the remaining more distant sample, or 2.8% in distance. The nearby sample may be fainter because it is not sampling the full distribution, given the small-number statistics. Alternatively, the lower signal-to-noise and poorer resolution of the more distant sample may be contributing a systematic uncertainty. (And both effects may also be at play.) However, it is well to keep in mind that at distances greater than 40 Mpc, the Cepheids are crowded by stars within an area 33 times larger than at 7 Mpc, while, simultaneously the signal-to-noise is lower for the more distant sample. (At distances greater than 20 Mpc, Cepheids are crowded by stars contained within an area at least 8 times larger than the equivalent resolution element (fitting radius) at 7 Mpc.) Future data will be required to assess the relative contributions of these two possibilities.
At this juncture in our CCHP program, we have a calibrating sample of 11 SNe Ia, which will continue to be augmented. For context, this sample can be compared with the early SHoES samples, where, for instance Riess et al. (2009) had a total of six nearby galaxies, for which they determined a value of = 74.2 3.6 km s-1 Mpc-1, a 4.9% measurement. Riess et al. (2011) then increased the sample by two additional SNe Ia, bringing their total to eight, and resulting in a value of = 73.8 2.4 km s-1 Mpc-1, a 3.3% uncertainty. The addition of these two supernovae increased the dispersion in the SNe Ia peak MB magnitudes from 0.08 to 0.12 mag and the mean absolute magnitude dropped significantly, from -19.34 mag down to -19.27 mag, a shift of 0.07 mag. On its own, this change would decrease the distance scale by 3.3% and increase the Hubble constant by a similar percentage. However, other improvements to the SHoES analysis led to a change in the opposite direction so that decreased by 0.8%. These results agreed extremely well with their current estimate of = 73.04 1.04 km s-1 Mpc-1, based on a sample of 42 SNe Ia.
Is the existing calibrating SNe Ia sample accurate enough to provide an estimate of at the 1% level? Some of the challenges in measuring accurate distances to the nearest galaxies in this sample can be seen in the lower panel of Figure 15. Shown is a comparison of the MB magnitudes for SNe Ia calibrator galaxies in common to the R16 and R22 studies (based on distances published in their Tables 5 and 6, respectively). The HST data for these original 19 galaxies (20 SNe Ia) were first analyzed in R16, and the same raw data were then re-analyzed in R22. Differences up to 0.3 magnitudes in distance modulus were measured, with a net increase in distance modulus of 0.033 0.024 mag, leading to an increase in MB (R22-R16) of 0.037 0.033 mag in the mean. Considered in isolation, this difference alone would result in a decrease in of 1.7% (but other improvements elsewhere resulted in no net change to ). However, most of the (most challenging distant) galaxies beyond 40 Mpc have not yet been imaged at the improved resolution of JWST.
The augmented R22 sample contains 42 SNe Ia, and its statistical uncertainty is smaller than available for our current JWST sample. However, it must be noted that 60% of the R22 sample galaxies are located at distances 20 Mpc and 25% are more than 40 Mpc distant. The more distant galaxies, on average, have lower signal to noise (S/N) and decreased resolution. Future high-resolution data will be needed to both improve the statistical precision of those distances and to rule out any further potential systematic effects.
In the left-hand panel of Figure 16, we compare the MB probability density functions for the SNe Ia in the R22 sample, those with combined SN Ia peak mB and Cepheid distance modulus errors ( ) less than 0.15 mag (23/42 objects). In blue are the 11 SNe Ia that are located closer than 31.8 mag in distance modulus (d 23 Mpc) ; in red are the 12 SNe Ia more distant than 31.8 mag. For this highest-quality (in both SNe Ia light curve and Cepheid distance precision) sample, there is again a clear systematic offset with distance in the peak SN Ia brightness, in the sense that the nearby sample is brighter in the mean. The right-hand panel shows the PDFs for the sample of SNe Ia with larger uncertainties ( 0.15). In this case, the peak brightness is fainter for both the nearby and more distant SNe Ia. These comparisons highlight once again the critical role that this intermediate rung is playing in the SNe Ia distance scale, and the need for additional JWST data and a larger SNe Ia calibrating sample.
11 Error Budget
The statistical and systematic uncertainties for each of the three distance indicators are discussed in detail in the companion papers of O24, H24 and L24. Here we note that in the determination of , the overall uncertainties for the individual galaxy distances become statistical uncertainties. For example, while the aperture corrections or reddening corrections for an individual galaxy contribute to its systematic uncertainty, for an ensemble of galaxies, the aperture corrections or reddening corrections become a source of random/statistical uncertainty. The overall statistical uncertainty is determined by the number of host galaxy calibrators, and the total number of SNe Ia contained therein.
In Table 5 we summarize the sources of the statistical and systematic uncertainties in our measurement. We list the statistical and systematic errors separately. The combined statistical errors are given by: . (The individual statistical uncertainties were discussed in §9.4.) The individual systematic errors, as discussed in detail in O24, H24, and L24 are 3.10, 1.54 and 1.90 km s-1 Mpc-1, for Cepheids, TRGB and the JAGB, respectively. The systematic errors common to all three methods are that for the zero point (anchored to NGC 4258), and the photometric zero point; otherwise the uncertainties are independent. A description of the components of the systematic errors is given in the notes to Table 5. Based on our Bayesian analysis with a flat prior, as described in §9.4 and illustrated in Figure 11, our adopted value of the Hubble constant and its uncertainty, as derived from our new JWST distances applied to the CSP sample of SNe Ia is = 69.96 1.05 (stat) 1.12 (sys) km s-1 Mpc-1.
Source of Error | Random Error | Systematic Error | Reference |
TRGB zero point | 2.5% | 2.2% aafootnotemark: | H24, §9.4 |
JAGB zero point | 2.7% | 2.8% bbfootnotemark: | L24, §9.4 |
Cepheid zero point | 2.6% | 4.3% ccfootnotemark: | O24, §9.4 |
Combined methods ddfootnotemark: | 1.5% | 1.6% | §9.4 |
11.1 Overall Systematic Uncertainties
Combining the results for the three different methods provides a means of averaging over the systematic effects that independently affect any given method. There are some systematics, however, that are common to all of the methods.
1.
As described in §10 and illustrated in Figures 15 and 16, a current uncertainty in the SN Ia distance scale is the relatively small number of calibrators with well-observed SNe Ia, and distances measured with Cepheid TRGB or JAGB distances. A concern is the trend that SN Ia absolute magnitudes get fainter with increasing distance (and increasing errors), a result that could be due to statistical effects and/or systematic effects in the measured distances, particularly in galaxies where crowding/blending effects could be more severe.
2.
The calibrations of the TRGB, JAGB stars and Cepheid Leavitt law for the JWST data presented in this paper are all based on the geometric distance to one galaxy, NGC 4258, and the three methods therefore share any systematic errors (known or unknown) in that determination. The currently cited total uncertainty in the NGC 4258 distance is 1.5% (Reid et al., 2019). Unfortunately, at 7.5 Mpc (and with a central accretion disk that is highly inclined, allowing accurate measurements of its maser orbits), NGC 4258 is the only nearby galaxy that is suitable for the calibration of the TRGB, JAGB stars and Cepheids based on masers. (The next nearest galaxy is at a distance of over 50 Mpc, with a measured uncertainty of 9% (Pesce et al., 2020).) With NGC 4258 alone, it is not possible to determine any underlying galaxy-to-galaxy differences inherent in the maser method, or to assess the impact of any systematics tied to assumptions in the modeling. In forthcoming papers, we will continue to improve upon and apply additional zero-point anchors to each of the three distance determination methods; for this study, we have adopted a consistent set of distances and calibration based on JWST data alone. We have explicitly tested the NGC 4258 calibration in the case of the TRGB (§12.3), finding excellent agreement with previous studies that were based on a calibration with four geometric anchors.
3.
The measurement of the Cepheid distance to NGC 4258 carries with it the uncertainty in , the slope of the sensitivity of the Leavitt law to metallicity, the value of which is still a subject of active debate in the literature. For the purposes of this study, we have adopted a value of = -0.2 mag/dex (as adopted by R22); this translates to less than a 1% (i.e., 0.7%) metallicity correction to , in the sense that decreases when the correction is applied.
However, if = -0.5 mag/dex as some Gaia studies have concluded (e.g., Ripepi et al., 2020), it would be a 2% correction. This becomes a more serious issue when using the LMC as an anchor galaxy. R22 adopt [O/H] = -0.29 for the LMC; and this would translate to a 6.3% metallicity correction for . Furthermore, the LMC’s geometric distance currently has the highest precision of all of the current anchors, making it the largest contributor to the estimation of in the multi-anchor analyses. Future work is needed to establish, with confidence, the metallicity sensitivity of the Leavitt law. In combining the three methods, an error in the Cepheid metallicity correction will enter as part of the overall systematic uncertainty in our value of .
The impact of metallicity on the Cepheid calibration presented here is small because the average metallicity of our sample is close to that of NGC 4258 itself, based on the metallicity scale inferred from HII regions adopted by R22. However, it is worth emphasizing that there remains a trade-off between having only a single anchor galaxy with a small metallicity correction versus the case of more anchor galaxies, each being impacted by increasingly uncertain metallicity corrections.
4.
For all three applications, we have adopted the same reddening law, which is assumed to be universal, with the same ratio of total-to-selective absorption, . In addition, for the JAGB and TRGB, where foreground reddening corrections are required (for the Cepheids a total reddening correction is measured), there is an uncertainty due to the zero-point calibration of the extinction law. Schlegel et al. (1998) estimates this uncertainty to be = 0.02 mag, which corresponds to an extinction uncertainty of AJ = 0.016 mag (0.7%). In an era of accurate cosmology, this level of uncertainty becomes more significant.
5.
In this analysis, we use SNe Ia to extend the distance range for the determination. Any remaining systematic error intrinsic to SNe Ia will be shared by all three of the methods presented here. These systematics might include inconsistent calibrations of photometric zero points across different surveys and instruments that can lead to systematic offsets (e.g., in the case of Pantheon+); in addition the calibrator light curves are generally not part of the well-calibrated CSP survey, which makes up the Hubble flow sample. Further uncertainties may arise in the corrections for dust extinction (e.g., existence of more than one type of dust: differences in the dust surrounding the SN Ia, variations in the extinction law, etc.), intrinsic scatter in the peak luminosity of SNe Ia, and/or effects such as the host galaxy mass step, which may not be fully constrained, potentially leading to residual, uncorrected biases in linking the calibrator SN Ia luminosities to the more distant SNe Ia.
6.
Finally, the results of an MCMC analysis may not be as robust as the confidence implied by the likelihood analysis. These include the fact that the probability might be accurately treated as a Gaussian near the peak, but it may not decrease as fast as a Gaussian in the tails; there may be unknown systematic errors revealed as the statistical uncertainty decreases and a systematic floor is reached; and the likelihood function may have uncertainties and/or assumptions that have not been included in the quoted uncertainty (e.g., Liddle, 2009; Hogg & Foreman-Mackey, 2018).
We note that the uncertainties that we are quoting are larger than those given, for example, by R22. In summary there are several reasons for this difference. (1) Our sample of SN Ia calibrators is smaller by a factor of four than that of R22, resulting in a factor of two larger statistical uncertainty. This uncertainty, however, is at least partially offset by the fact that we have three distance methods. (2) We have only a single anchor galaxy, NGC 4258, whereas R22 have three anchors (LMC, the Milky Way and NGC 4258). However, as discussed in §12.3, there is excellent agreement with the -band calibration of the TRGB with four anchors. (3) We have adopted a more conservative estimate of the uncertainty in the metallicity dependence of the Leavitt law, reflecting the current uncertainty in the Gaia absolute calibration, as well as the range in published values of the metallicity slope parameter, (see §11.1). (4) We separate our statistical and systematic uncertainties and do not marginalize over the uncertainties.
12 Comparisons With Previously Published Data
12.1 TRGB Comparison with Archival SHoES Data for NGC 4258
Archival JWST data from the SHoES Cycle 1 program (GO-1685 , PI: A. Riess) offers an important constraint on our TRGB calibration, taken with different filters than our own data. In the left panel of Figure 17, we show a versus () color-magnitude diagram for a halo region of the galaxy NGC 4258 based on archival JWST data from the SHoES team. The position of the TRGB is shown by the solid black horizontal line at F090W = 25.08 mag. The right panel shows the marginalized and smoothed luminosity function in black, and the Sobel edge-detection filter output is shown in blue. Detection of the tip in the left panel, and its quantitative measurement in the right panel, is unambiguous. The luminosity function was smoothed by 0.1 mag before its digital first derivative was output by the Sobel filter.
The JWST -band () data were derived from the sum of four 258-second NIRCam exposures; the four equivalent exposures were taken in parallel, directed to the beam-split long-wavelength channel of NIRCam. The Sobel response function has a measured width of 0.10 mag, within which there are 624 red giant branch stars contributing to the measurement. We conservatively adopt an uncertainty of 0.02 mag for the statistical uncertainty. A foreground -band galactic line-of-sight extinction of 0.02 mag, to the halo of NGC 4258, as derived by Anand et al. (2024a) is adopted here. This gives a reddening-corrected magnitude of 25.06 mag for the TRGB in NGC 4258. Adopting a geometric distance modulus of 29.397 0.032 mag (Reid et al., 2019) yields our zero-point calibration of MF090W = -4.336 0.02 (stat) 0.032 [sys] mag. This can be compared to the zero point of MF090W = -4.32 0.025 (stat) mag published by Newman et al. (2024), averaging over six nearby galaxies observed with JWST, using the same filter combinations as discussed here, but zeroed to previously published TRGB distances using an -band calibration of MI = -4.05 mag (Freedman, 2021). The SHoES team (Anand et al., 2024b) has also published an averaged zero point for the TRGB method giving MF090W = -4.362 0.033 (stat) 0.045 [sys] mag (with different applications of the method differing by up to 0.04 mag). In all cases, the zero-point agreement is very good: Anand et al. is -0.025 mag brighter than our calibration, while Newman et al. (2024) are +0.017 mag fainter. Taking the average of the three determinations would return our zero point to within 0.003 mag.
12.2 Comparison of Previously Published TRGB and JAGB Distances
In Figure 18, we compare measured TRGB distances with those from JAGB stars, including nearby galaxies measured with ground-based telescopes, and tabulated in Madore & Freedman (2024a, b); Lee et al. (2024b). There is strong consistency, with remarkably low scatter amongst these two distance scales, over a range of a factor of 50 in distance. These comparisons provide an additional external constraint on both the zero-point calibrations of the TRGB and JAGB methods, as well as their respective internal precisions.
12.3 Comparison of H0 Values: -band TRGB Distances Measured Using HST
In Freedman et al. (2019, 2020); Freedman (2021) we presented results from a HST program to measure -band TRGB distances based on observations of 20 SNe Ia located in a sample of nearby galaxies. In Freedman (2021) four geometric anchor distances were applied to the band TRGB distance scale: (1) Gaia parallaxes for Milky Way globular clusters (Cerny et al., 2020; Maíz Apellániz et al., 2021; Vasiliev & Baumgardt, 2021); (2) the detached eclipsing binary distance (DEB) to the LMC (Hoyt, 2023; Pietrzyński, 2019); (3) the DEB distance to the SMC (Hoyt, 2023; Graczyk et al., 2020) and (4) the maser distance to NGC 4258 (Reid et al., 2019). This calibration, based on four geometric anchors, resulted in a value of = 69.8 0.6 (stat) 1.6 [sys] km s-1 Mpc-1, a number that is in excellent statistical agreement with the JWST analysis presented in this current study.
Of the 20 SNe Ia in Freedman (2021) and the 10 galaxies in the current study with TRGB distances, there are six galaxies in common to both studies: M101, NGC 1365, NGC 4038, NGC 4424, NGC 4536 and NGC 5643. The differences in distance moduli (in the sense JWST -band minus HST -band) are +0.098, -0.028, -0.070, -0.073, -0.026, and +0.137, respectively, with a weighted average distance modulus offset amounting to only +0.017 mag 0.014 (error on the mean) mag.
We note that the data for these two studies are entirely independent: the -band data for the target galaxies were obtained with HST, whereas the -band data were obtained with JWST. Furthermore, and importantly, the -band distances were measured with respect to the LMC, the SMC, the Milky Way and NGC 4258, whereas the -band distances were measured with respect to NGC 4258 alone. This independent comparison serves to provide a strong external check on the TRGB distance scale.
12.4 H0 With Time
While both the precision and the accuracy of the measurements of distances to galaxies have improved considerably in recent years, there are still reasons to remain open as to whether our currently estimated uncertainties truly reflect the total uncertainties. Shown in Figure 19 are all published values from the database maintained by Ian Steer (Steer (2020), Steer (2024, priv. comm.) (small dark blue dots) and the values of calibrated by Cepheids (larger red filled circles).
There are several points worth noting in Figure 19. First, the values of obtained from SNe Ia, as calibrated by Cepheids, have remained remarkably constant over a quarter of a century. Second, the Cepheid-based values of are systematically higher than the (mean, median or mode) of the distribution of other determinations (Steer et al. 2020, 2024). Third, there is no bimodal distribution of published values centered, respectively, around the oft-quoted values of 67 and 73 km/s/Mpc. Future studies will be required to unambiguously confirm the higher values that are based on Cepheid measurements alone or determine whether systematic errors ultimately prove to be the explanation.
13 Future Prospects
With JWST, we have the means of improving both the accuracy and the precision of the locally determined extragalactic distance scale, solidly providing a multiply calibrated and independently verified determination of current expansion rate of the universe, .
1.
Improving the Zero-Point (Anchor) Calibration: There are four galaxies at present for which geometric distances can be measured for calibrating astrophysical/stellar distance indicators suitable for determining the extragalactic distance scale. These include the LMC, NGC 4258, and the SMC, with quoted uncertainties of 1% (Pietrzyński, 2019), 1.5% (Reid et al., 2019), 2% (Graczyk et al., 2020), respectively, and an estimated range of 1-3%. The Milky Way parallax offset currently has a significant impact on derived parallaxes for Cepheids and the period-luminosity-metallicity relation (e.g., see Groenewegen, 2021; Owens et al., 2022, for a discussion of the current uncertainties). Future data releases for the Gaia satellite are forecast to provide a 1% parallax calibration for the Cepheid PL relation. In addition, accurate parallax measurements for Milky Way globular clusters will provide a zero-point calibration for the TRGB, again at a level of 1%.
2.
Increasing the Numbers of SN Ia Host Calibrators: On average, only about one new SN Ia, per year, is found within a distance accessible for follow-up discovery of Cepheids, TRGB or JAGB stars using HST. Thus the total sample of SN Ia host galaxies is not expected to increase significantly over the next decade. At present there are 37 galaxies for which HST Cepheid distances have been measured as part of the SHoES project. Ten supernova-host galaxies have been observed with JWST as part of this paper (GO-1995), and an additional 11 have been observed with JWST for programs GO-1685 and 2875. One galaxy is in common to both programs, NGC 5643, and both programs additionally have observations of NGC 4258. JWST GO-3055 includes observations in the outer regions of NGC 1404 and NGC 1380 with the filters and , optimal for TRGB and JAGB measurements. In the case of JWST, doubling the sample of galaxies observed will improve the precision of by . Beyond improving the statistical uncertainties alone, observations of more distant galaxies where crowding effects are more severe, will be important for constraining potential systematic uncertainties increasingly encountered as a function of distance.
Currently planned 30-meter-class ground-based optical telescopes (e.g., the Giant Magellan Telescope) will have 10 times the resolution of HST, and will allow Cepheids, TRGB and JAGB stars to be discovered and measured within a 1,000 times greater volume, thereby providing a significant increase in the numbers of SN Ia host galaxies. The Cepheids can then be followed up in the near-infrared with JWST or the Nancy Grace Roman Space Telescope (see below).
3.
Photometrically Consistent and Wider Areal Coverage of Nearby Galaxies: The Nancy Grace Roman Space Telescope888https://roman.gsfc.nasa.gov/, due to be launched in 2027, will have 100 times the field of view (FOV) of HST, with comparable resolution and sensitivity, with wavelength coverage extending from 0.48 to 2.3 m. The large FOV will enable the mapping of entire disks and halos of nearby SNe Ia host galaxies, giving simultaneous measurements of the TRGB, Cepheids, and JAGB stars in single (massive) pointings. The simultaneous measurements will provide a consistent photometric calibration for all three methods.
4.
Future SN Surveys: Over the last few decades, an enormous amount of time and energy has gone into ground-based surveys for SNe Ia (e.g., Pantheon+). This database is the merger of observations from 18 different surveys (Scolnic et al., 2022), taken with different telescopes, instruments, calibrations, etc. The Vera C. Rubin Observatory999https://rubinobservatory.org/, which will begin taking data in 2025, is forecast to discover over 3 million supernovae in its first decade of operation. Follow-up spectroscopy to measure redshifts using other ground-based facilities will pave the way for a dataset that has a consistent photometric calibration, without the need to apply photometric offsets and/or corrections to heterogeneous samples, as is the case today.
5.
Additional Tests for Systematics: Importantly, future measurement of employing techniques completely independent of the local distance scale, accurate at the 1-2% level, will be essential for ruling out remaining systematic uncertainties in the local distance scale. An example is gravitational wave sirens (e.g., Chen et al., 2018), which, if more objects can be discovered, hold exciting promise.
6.
More Accurate Modeling to Standardize SNe Ia: A promising path to increasing precision and accuracy in , at once addressing Points 2 and 5, is improvements to models for standardizing SN Ia Hubble residuals (e.g. Boone et al., 2021; Stein et al., 2022). For these models, the SN Ia Hubble residuals can be standardized to 0.07-0.08 mag in distance using time-series spectrophotometry (compared to typical dispersions in the 0.15 mag range for light curve standardization approaches). The smaller intrinsic dispersion of such models can both significantly increase the amount by which statistical precision of tightens per SN Ia calibrator and, perhaps more importantly, also tighten the constraints on any as-yet unseen systematics in SN Ia standardization, which are more likely to disproportionately impact the measurement of due to the currently small number of calibrator SNe Ia.
14 Summary and Conclusions
The infrared sensitivity and high resolution of the JWST is providing a powerful new means of measuring the distances to nearby galaxies, and thereby enabling new and independent determinations of . In this paper, we have measured the distances to 10 nearby galaxies using three independent astrophysical distance indicators: Cepheids, the TRGB and JAGB/carbon stars. SNe Ia have previously been well observed in all of these galaxies. For the JAGB distance scale, the data analysis from the raw data frames through to determination of was carried out blind.
An inter-comparison of the galaxy-to-galaxy distances results in agreement with a combined scatter in each case (i.e., Cepheid versus TRGB, TRGB versus JAGB, JAGB versus Cepheid) of less than 4%. This agreement represents a remarkable improvement from recent decades. In the case of the JAGB and TRGB distances, the results are even more striking, with agreement at a level of less than 1%.
A significant result from the current study is that the two distance indicators (TRGB and JAGB) that are least affected by crowding/blending or reddening, and that are based on a single, consistent calibration from JWST/ NIRCam, are giving larger distances (and therefore result in a lower ) than for the Cepheids. This difference in measurement of the local distances is independent of the ultimate step taken in applying the distances to SNe Ia, and thereby determining in the unperturbed cosmic flow. These differences are pointing to systematics affecting one or more of the distances and need to be better understood. However, while they do not rule it out, the results presented here do not lend strong support to the suggestion that there is missing fundamental physics in the early universe. Only future data will settle this issue unambiguously. Currently, the distances obtained using the TRGB and JAGB methods agree extremely well, and these methods have the advantage that the inner disks of galaxies can be avoided in their application, unlike Cepheids, which are young and embedded in regions of dust and high source densities and high surface brightness.
One of the largest remaining uncertainties in measuring the local value of using SNe Ia (e.g., SHoES or CCHP) is the small number of nearby, calibrating SN Ia host galaxies. This small sample is a consequence of the fact that SNe Ia are relatively rare, and as a result, there are few galaxies in the local volume containing SNe Ia that are also close enough to resolve Cepheids, TRGB or JAGB stars with HST. Moreover, these serendipitously available supernovae also appear to be amongst the intrinsically brightest. But, in being nearby and well observed, they also have the lowest uncertainties. More data will be required to ascertain if there is a systematic error in the distances to the more distant SN Ia galaxies. In addition, the sample needs to be enlarged to simultaneously address potential bias and decrease the statistical uncertainties. Increasing this sample will require more time; unfortunately only 1-2 of these SNe Ia occur each year. In the context of resolving the tension, it is critical to decrease both the systematic and statistical uncertainties.
Finally, we summarize our results in Figure 20. It shows the good agreement between the two cosmological values from Planck (Planck Collaboration et al., 2020), DESI (DESI Collaboration et al., 2024), and the three local distance scale values (and their averaged value) from CCHP JWST. The accuracy of our measurements is dominated by the systematic uncertainty in the distance to the anchor galaxy, NGC 4258, which is currently estimated to be 1.5%. Based on our three methods, applied to the CSP sample of SNe Ia, our best estimate of is 69.96 1.05 1.12 km s-1 Mpc-1. Individually, the values for the three methods agree well, with only 4 km s-1 Mpc-1 separating the largest and smallest values. The value of from this study falls between that obtained from studies of the CMB, from BAO measurements calibrated by BBN, and the SHoES value, all without significant tension. Future work on the local distance scale will shed light on whether or not additional physics is needed beyond the CDM cosmological model.